Design an algorithm to solve the problem using pseudocode:



(Print a table) Write a program that displays the following table:



p p*5 p*10



5 25 50


10 50 100


25 125 250


50 250 500

1
Expert's answer
2022-01-05T08:13:38-0500
Start
   Declare variable p
   Declare variable tempP
   Input p
   tempP=p
   Display p," ", p*5," ",p*10
   p=tempP
   p=p*2
   Display p," ", p*5," ",p*10
   p=tempP
   p=p*5
   Display p," ", p*5," ",p*10
   p=tempP
   p=p*10
   Display p," ", p*5," ",p*10
Stop